# Perceptual Decoupling for Scalable Multi-modal Reasoning via Reward-Optimized Captioning

### v1

@@ Abstract
Recent advances in slow-thinking language models (e.g., OpenAI-o1 and DeepSeek-R1) have demonstrated remarkable abilities in complex reasoning tasks by emulating human-like reflective cognition. However, extending such capabilities to the multi-modal large language models (MLLMs) remains challenging due to the high cost of retraining vision-language alignments when upgrading the underlying reasoner LLMs. A straightforward solution is to decouple perception from reasoning, i.e., converting visual inputs into language representations (e.g., captions) that are then passed to a powerful text-only reasoner. However, this decoupling introduces a critical challenge: the visual extractor must generate descriptions that are both faithful to the image and informative enough to support accurate downstream reasoning. To address this, we propose Reasoning-Aligned Perceptual Decoupling via Caption Reward Optimization (RACRO) —a reasoning-guided reinforcement learning strategy that aligns the extractor’s captioning behavior with the reasoning objective. By closing the perception-reasoning loop via reward-based optimization, RACRO significantly enhances visual grounding and extracts reasoning-optimized representations. Experiments on multi-modal math and science benchmarks show that the proposed RACRO method achieves state-of-the-art average performance while enabling superior scalability and plug-and-play adaptation to more advanced reasoning LLMs without the necessity for costly multi-modal re-alignment.
@@ 1 Introduction
Recent developments in slow-thinking language models, such as OpenAI-o1 [26], DeepSeek-R1 [22], Gemini-Thinking [11], and Qwen3 [45], have led to significant advancements in the performance of large language models (LLMs) for complex math and science tasks. These models replicate human-like reflective thinking, intentionally modeling reasoning as a step-by-step process rather than relying on superficial shortcuts. As a result, they surpass fast-thinking models like GPT-4o [25] and Claude 3.5 [2], showing improvements of over 30% on mathematical benchmarks such as AIME24 [1] and AMC23 [15], along with approximately 10% gains on science benchmarks like GPQA [41].
Motivated by these successes, there is growing interest in equipping multi-modal large language models (MLLMs) with similar slow-thinking capabilities to solve vision-based reasoning tasks that require both spatial perception and logical inference. A common strategy—adopted by methods like VL-Rethinker [47] and MM-EUREKA [38]—is to fine-tune a pre-trained MLLM (e.g., Qwen2.5-VL [5]) on multi-modal reasoning datasets using reinforcement learning with verifiable reward signals. However, this approach faces two critical limitations: (1) it is inherently restricted by the capabilities of underlying LLMs used for initialization; (2) adapting to more advanced LLMs necessitates redoing the vision-language alignment process, which involves burdensome multi-modal training that requires computing trillions of tokens. This raises a challenging research question: Can we efficiently integrate slow-thinking strategies into multi-modal reasoning without retraining the entire multi-modal LLM?
A seemingly intuitive solution is to utilize a perceptual decoupling approach: separate the vision and reasoning components by converting multi-modal inputs into natural language captions, which are then passed to an external, more powerful text-only reasoner. While this idea promises flexibility and modularity, it introduces a fundamental challenge—the success of the reasoning process now critically depends on the quality of intermediate captions. If the visual extractor generates captions that are incomplete, overly generic, or hallucinatory, even the strongest reasoning LLMs will fail to give correct reasoning processes. As shown in Figure 3, existing MLLMs might struggle to produce reasoning-relevant descriptions, especially for complex mathematical and scientific visual inputs [51].
Figure 1: Scalable multi-modal reasoning of RACRO. RACRO reveals better scalability compared to the traditional visual alignment, while remaining computationally efficient.
To overcome this limitation, we propose Reasoning-Aligned Perceptual Decoupling via Caption Reward Optimization (RACRO) —a unified framework that not only decouples visual perception from reasoning, but also ensures their alignment through a reasoning-aware reinforcement learning procedure. In the proposed RACRO method, an MLLM-based extractor is responsible for encoding task-relevant visual information to natural language, while an external text-only LLM reasoner handles the slow-thinking reasoning process. Crucially, we introduce Caption Reward Optimization (CRO), a reinforcement learning algorithm that fine-tunes the extractor using reasoning outcomes as feedback. Given an input image and a question, the extractor generates multiple caption candidates; each is passed to the reasoner, and rewards are assigned based on whether the reasoner generates correct answers or not. This feedback loop aligns the extractor with the reasoning objective, guiding it to generate faithful and problem-relevant descriptions.
By integrating perceptual decoupling with CRO into a unified framework, we enable effective and scalable slow-thinking in multi-modal reasoning systems. Experiments on challenging benchmarks—such as MathVerse [56], MathVision [48] and LogicVista [52]—show that the proposed RACRO method achieves state-of-the-art results. Furthermore, the proposed RACRO method supports plug-and-play adaptation to novel reasoning LLMs without extra multi-modal retraining, making it a practical and scalable solution for the rapid evolution of foundation models. It is illustrated in Figure 1, comparing the average performance of various MLLMs and their RACRO-enhanced variants across model sizes. For each RACRO-enhanced model group (e.g., Qwen2.5-VL-3B ($\bf{RACRO}$)), the original variant (leftmost, solid) is only trained with a base LLM reasoner with a small amount of data (e.g., 39K) using CRO, which can then achieve scalable performance with stronger LLMs (see detailed configurations on the choices of LLMs in Appendix A). Connecting these points yields a novel Pareto front that captures a superior trade-off between model size and performance than other base MLLMs.
Our contributions can be summarized as the following three perspectives:
  • We propose RACRO, a unified framework that decouples visual perception from reasoning and aligns them through reinforcement learning with verifiable rewards (RLVR).
  • We introduce Caption Reward Optimization, a novel reinforcement learning strategy that uses reasoning success as a reward signal to optimize the extractor’s captioning behavior.
  • We obtain state-of-the-art performance on multi-modal reasoning benchmarks, while maintaining scalability and eliminating the necessity for costly vision-language re-alignment.
(a) Existing MLLM alignment methods
(b) RACRO (ours)
@@
Figure 2: Comparisons between existing alignment methods for reasoning MLLMs and RACRO. For novel LLMs, existing methods (a) repeatedly conduct the intensive alignment process, while (b) RACRO decouples visual perception from reasoning (Sec. 3.1) by learning to extract reasoning-aligned visual contexts with the proposed CRO paradigm (Sec. 3.2).
@@ 2 Related Work
Multi-modal Large Language Models (MLLM).
With the development of LLMs, integrating visual perception capability into LLMs has received increasing attention [19, 27, 30]. To address the gap between different modalities, existing paradigms [5, 20, 21, 59] start with a text-only LLM and conduct pre-training on large-scale image captioning data to obtain a vision-language connector. Then, supervised fine-tuning (SFT) is applied with visual instruction tuning datasets. It is important to note that this procedure is computationally intensive. For example, Qwen2.5-VL [5] consumes around 4.1T tokens, and Kimi-VL [44], a more recent MLLM with MoE architecture, requires 2.3T tokens for the alignment. Under this paradigm, it remains formidable to upgrade MLLMs to newer LLMs. Instead, this paper proposes a computationally efficient solution for multi-modal reasoning.
Empowering MLLMs with Slow-Thinking Capabilities.
To empower MLLM with similar slow-thinking abilities, a widely adopted approach is to start with an existing MLLM (e.g., Qwen2.5-VL [5]) and perform reinforcement learning or knowledge distillation. For example, VL-Rethinker [47], MM-EUREKA [38] and NoisyRollout [31] apply GRPO [42] (or its variants [13, 31, 32, 38]) to MLLMs to learn deliberate reasoning patterns. Distillation-based methods, such as the OpenVLThinker [13], R1-OneVision [54], and Vision-R1 [24], perform supervised fine-tuning (SFT) on the vision-language reasoning data generated via a caption-and-reasoning pipeline [19], where visual inputs are converted to textual descriptions first, and strong text-only LLMs are then utilized to generate reasoning traces. However, their performances are limited by the base LLM (e.g., Qwen2.5 [53]), which often lags behind state-of-the-art reasoning models (e.g., DeepSeek-R1-Distilled [22]) in tasks like math. While replacing it with a stronger LLM is an intuitive solution, re-aligning vision and language through full retraining on trillions of tokens is prohibitively costly. In this work, we introduce the RACRO, a reasoning-aligned perceptual decoupling framework that bridges perception and reasoning via reward optimization, allowing seamless integration of stronger LLMs without retraining the vision-language interface, and meanwhile, demonstrating superior scalability w.r.t model sizes (c.f., Figure 1).
@@ 3 Methodology
In this section, we propose Reasoning-Aligned Caption Reward Optimization (RACRO), a unified framework that enables multi-modal slow-thinking reasoning by aligning the visual perception module with the downstream reasoning objectives. RACRO builds on the intuition of perceptual decoupling, where visual inputs are first translated into natural language representations that can be processed by powerful text-only reasoners. However, rather than treating the decoupling as static and assuming fixed captions suffice, RACRO closes the perception-reasoning loop through reinforcement learning, optimizing the captioning behavior of the visual extractor for downstream reasoning success.
3.1 Reasoning-Aligned Perceptual Decoupling
The perceptual decoupling paradigm is based on the assumption that the process of solving visual reasoning tasks can be separated into two distinct stages: visual perception and text-only reasoning. Specifically, given an image and a related question, if all relevant visual details can be losslessly converted into natural language, the problem can be solved using a text-only LLM. This decoupling allows for seamless upgrades to more powerful LLMs without the need to retrain or realign vision-language models. However, extracting all necessary visual details remains a severe challenge, as visual information can be dense, complex, and ambiguous [19, 51]. To address that, we propose the Reasoning-Aligned Perceptual Decoupling framework, focusing on providing reasoning-aware visual details to the LLM instead of attempting to capture the entire visual input. Figure 2 gives an overview of the framework. Specifically, It consists of two steps:
Figure 3: Inaccurate visual captions result in reasoning failure. MLLMs might produce hallucinated (left), vague, or incomplete captions (middle), particularly for complex visual reasoning, while captions containing necessary query-relevant contexts are essential for successful reasoning (right).
Extraction.
Guided by the goal of reasoning-aligned perceptual decoupling, we design an extraction process that generates visual details tailored to the reasoning needs for a given question. Instead of producing generic descriptions, we leverage an MLLM, such as Qwen2.5-VL [5] or InternVL3 [59], denoted as $M_{V}$, to produce two complementary outputs from an image $I$ and a query $\mathbf{q}$: a query-conditioned caption $\mathbf{c}=M_{V}(I,P_{\text{cap}}(\mathbf{q}))$, and a tentative solution $\hat{\mathbf{y}}=M_{V}(I,P_{\text{cot}}(\mathbf{q}))$, with each generated using distinct prompts (see Appendix B). The caption $\mathbf{c}$ captures visual content which is selectively relevant to the question, functioning as a reasoning-aligned abstraction. Although $\hat{\mathbf{y}}$ may not be fully accurate due to the limited reasoning capabilities of MLLMs, it often contains useful inductive signals that can inform and guide further downstream reasoning inference.
Reasoning.
The reasoning component is separately handled by a text-only large language model (LLM) (e.g., R1-Distilled-7B/32B [22], Qwen3-8B [45], or QwQ-32B [46]), denoted as $M_{R}$. It takes as input the question $\mathbf{q}$, the reasoning-aligned caption $\mathbf{c}$, and the tentative solution $\hat{\mathbf{y}}$, and produces the final answer: $\mathbf{y}=M_{R}(P_{\text{reason}}(\mathbf{q},\mathbf{c},\hat{\mathbf{y}}))$, where $P_{\text{reason}}$ is a prompt formatting function that integrates the inputs into a coherent textual reasoning context (see Appendix B).
A core advantage of this decoupled framework is its flexibility: by separating reasoning from visual perception, we can boost performance on multi-modal tasks simply by upgrading to more capable LLMs for $M_{R}$, without modifying or retraining the MLLMs. While concurrent works have adopted similar decoupling paradigms—either as data engines for generating reasoning traces [24, 54] or as the inference framework [37]—they fail to investigate how to generate 1) query-conditioned and 2) reasoning-aligned visual representations, which is the focus of this paper in the following.
3.2 Caption Reward Optimization
While this design enables flexible upgrade of the reasoner without vision-language re-alignment, it exposes the system to a critical bottleneck: the quality of the caption directly affects the reasoning outcome. In practice, existing MLLMs may generate hallucinated [51], vague, or incomplete captions, particularly for complex visual reasoning tasks such as the geometric math questions (see Figure 3). To address this bottleneck, RACRO introduces Caption Reward Optimization—a reasoning-guided reinforcement learning algorithm that fine-tunes the captioning extractor using the feedback from a text-only reasoner, aiming at reinforcing captions that maximize likelihood of successful reasoning.
Caption Reward Optimization

### v2

# Reasoning-Aligned Perception Decoupling for Scalable Multi-modal Reasoning

@@ Abstract
Recent breakthroughs in reasoning language models have significantly advanced text-based reasoning. On the other hand, Multi-modal Large Language Models (MLLMs) still lag behind, hindered by their outdated internal LLMs. Upgrading these is often prohibitively expensive, as it requires complete vision-language alignment retraining which is costly. To address this issue, we introduce Perception-Reasoning Decoupling, which modularizes the MLLM’s reasoning component and makes it easily replaceable. This approach redefines the MLLM’s role to convert multi-modal inputs into detailed textual outputs that can be processed by any powerful, external, text-only LLM reasoners. To align the MLLM’s perceptual output with the final reasoning task, we propose a novel reinforcement learning algorithm called Visual Perception Optimization (VPO). VPO rewards the MLLM based on the correctness of answers generated by the external reasoner to produce faithful and query-relevant captions. Together, this decoupling pipeline and VPO form our Reasoning-Aligned PerceptIon Decoupling (RAPID) approach. Empirical results show that RAPID achieves significant performance gains on multi-modal reasoning benchmarks. Crucially, RAPID enables a novel inference-time scaling paradigm: Once trained with VPO, the MLLM can be paired with any state-of-the-art LLM reasoner for consistent performance improvement without retraining.
@@ 1 Introduction
Recent reasoning language models, such as OpenAI-o1 (Jaech et al., 2024) and Qwen3 (Yang et al., 2025a), have driven significant gains in complex math and science tasks. By emulating a deliberate, step-by-step reasoning process akin to human reflection, these models avoid superficial shortcuts. As a result, they substantially outperform previous models like GPT-4o (Hurst et al., 2024), with improvements exceeding 30% on math benchmarks like AIME24 (AIME, 2024) and around 10% on science benchmarks like GPQA (Rein et al., 2024).
Translating breakthroughs from the uni-modal text to the multi-modal domain remains a significant challenge. Existing multi-modal large language models (MLLMs), like Qwen2.5-VL (Bai et al., 2025), Gemma3 (Team et al., 2025a), and InternVL3 (Zhu et al., 2025), still struggle with reasoning and math-intensive tasks because their underlying LLMs are outdated or lack slow-thinking capabilities. While approaches like VL-Rethinker (Wang et al., 2025a) and MM-EUREKA (Meng et al., 2025) try to improve performance with reinforcement learning, their success is fundamentally restricted by the reasoning capability of the base LLM. The ideal solution, namely, switching the LLM with the most state-of-the-art one, is often prohibitive, as it requires repeating the entire, costly vision-language alignment process. This raises the critical question: Can we replace the LLM within an MLLM to unlock advanced reasoning [^1] In this paper, we focus on multi-modal math and science reasoning tasks. efficiently, without undertaking redundant vision-language retraining?
To address that, we propose the Perception-Reasoning Decoupling pipeline, where we re-focus the MLLM’s primary role on perception. It first translates the multi-modal inputs into a comprehensive textual representation, which is then processed by a separate, powerful, external LLM for reasoning. This decoupling allows flexible alteration of the LLM reasoner, offering a path to circumvent the costly retraining cycle. Our key distinction from similar two-stage pipelines (Tiong et al., 2022; Guo et al., 2022; Hu et al., 2022; Gou et al., 2024; Lu et al., 2025) lies in the textual representation, which includes both a query-relevant caption and a tentative solution to ensure all essential visual information is captured for subsequent reasoning. However, the critical challenge in this new pipeline is that the generated textual outputs are not optimized for correct reasoning.
Figure 1: Comparisons on multi-modal reasoning benchmarks on average performance and total model size between RAPID-enhanced Qwen2.5-VL series of models and the other existing MLLMs. Check the detailed numerical results in Appendix A and experimental settings in Sec. 4.1.
To overcome that, we introduce Visual Perception Optimization (VPO), a novel policy gradient algorithm operating through a reinforcement learning feedback loop where, given a user query, the MLLM first generates a group of query-relevant captions, which are then considered as contexts for the external LLM reasoner to generate final answers. With a correctness-based reward function, the perceptual MLLM is aligned with the reasoning objective, and guided to generate faithful and query-relevant captions optimized for the correctness of downstream reasoning. The combination of the Perception-Reasoning Decoupling pipeline with the VPO algorithm forms our overall approach, named Reasoning-Aligned PerceptIon Decoupling (RAPID)
Empirically, RAPID achieves notable performance gains on challenging benchmarks such as MathVerse (Zhang et al., 2024), MathVision (Wang et al., 2024b) and LogicVista (Xiao et al., 2024). Moreover, as perception and reasoning are decoupled, the MLLMs trained with VPO generate textual outputs that can be directly fed to any LLM for reasoning. This eliminates the necessity for retraining, and enables RAPID to be a practical solution for the rapid evolution of MLLMs and reasoning LLMs. Figure 1 compares various MLLMs against the RAPID-enhanced Qwen2.5-VL series. For each RAPID-enhanced group (e.g., Qwen2.5-VL-3B), we optimize the MLLM with VPO using minimal data (39K). The resulting performance curves are generated simply by pairing the optimized MLLM with increasingly powerful external LLMs (see Appendix A for the choice of configurations), demonstrating a novel inference-time scaling paradigm.
Our contributions can be summarized as follows:
  • We introduce the Perception-Reasoning Decoupling pipeline, which redefines MLLMs’ focus to multi-modal perception, allowing the reasoning component to be flexibly replaced by any advanced external LLM without burdensome retraining.
  • We propose Visual Perception Optimization (VPO), a novel policy gradient algorithm that aligns the MLLM’s perceptual outputs by using the correctness of the external LLM’s final answers with the perceptual outputs as contexts for reward signals.
  • Combining both, RAPID achieves significant performance gains and introduces an efficient, novel “plug-and-play” inference-time scaling approach. By eliminating the costly retraining required by traditional methods, an one-time optimized MLLM can be paired with any stronger LLM for continual performance improvements, as demonstrated in Figure 1.
(a) Existing MLLM alignment methods.
(b) RAPID (ours).
@@
Figure 2: Comparisons between RAPID and existing alignment methods for reasoning MLLMs. For novel LLMs, existing methods (a) repeatedly conduct the compute-intensive alignment procedure, while (b) RAPID decouples the visual perception from text-only reasoning (Sec. 3.1) by learning to extract reasoning-aligned visual contexts with the proposed VPO algorithm (Sec. 3.2). Note that the caption penalty, as in Eq. 3, is omitted here for simplicity. The flame and snowflake icons indicate the models are trainable and frozen, respectively, during the process.
@@ 2 Related Work
Improving the Reasoning Ability of MLLMs.
Start with an existing MLLM (e.g., Qwen2.5-VL (Bai et al., 2025)), a widely adopted approach is to perform reinforcement learning or knowledge distillation. For example, VL-Rethinker (Wang et al., 2025a), MM-EUREKA (Meng et al., 2025) and NoisyRollout (Liu et al., 2025a) apply GRPO (Shao et al., 2024) (or its variant (Liu et al., 2025b)) to MLLMs to learn deliberate reasoning patterns. Distillation-based methods, such as R1-OneVision (Yang et al., 2025b), Vision-R1 (Huang et al., 2025) and ReVisual-R1 (Chen et al., 2025) perform supervised fine-tuning (SFT) on reasoning data. However, both methods are restricted by the base LLMs (e.g., Qwen2.5 (Yang et al., 2024)), which lags behind state-of-the-art reasoning models (e.g., Qwen3-8B (Yang et al., 2025a)). While adopting a stronger LLM is an intuitive solution, re-aligning vision and language through full retraining on trillions of tokens is prohibitively costly.
Caption-then-Reason Pipelines.
To leverage LLM reasoning without intensive retraining, prior work explores similar “caption-then-reason" pipelines that decouple perception from reasoning. These approaches use Vision-Language Models (VLMs) (Radford et al., 2021; Li et al., 2022; Yu et al., 2022) or MLLMs for the perception task, while a separate LLM handles reasoning. Efforts in this area primarily focus on improving caption generation, for instance by selecting query-relevant image patches for captioning (Tiong et al., 2022; Guo et al., 2022), prompting MLLMs for query-aware captions (Gou et al., 2024), or enhancing captioning datasets (Hu et al., 2022; Lu et al., 2025). RAPID differs from these works in two key aspects. First, it includes a tentative solution in its generated output to better capture critical visual information. Second, while existing methods do not optimize captions for the final outcome, RAPID rewards the captioning process based on the correctness of the final answer produced by the reasoning LLM.
@@ 3 Methodology
This section describes the two main components of RAPID: perception-reasoning decoupling (Sec. 3.1) and visual perception optimization (Sec. 3.2). Figure 2 gives an overview of the approach.
3.1 Perception-Reasoning Decoupling
Given an image $I$ and a relevant query $q$, our perception-reasoning decoupling pipeline involves two consecutive stages: (i) Perception, where an MLLM (e.g., Qwen2.5-VL (Bai et al., 2025)) acts as a perception module to generate a group of textual outputs $O_{p}$ (detailed below) with respect to the image $I$ and a perception prompt. (ii) Reasoning, where a powerful text-only LLM reaonser (e.g., R1-Distilled-7B (Guo et al., 2025) or Qwen3-8B (Yang et al., 2025a)) receives the original query $q$ and a consolidated set of perceptual outputs, $O_{p}$, which are structured by a reasoning prompt $P_{r}$ (shown in Fig. 12): $y=\operatorname{LLM}(P_{r}(q,O_{p}))$. A key advantage of this decoupling pipeline is that the textural outputs form a universal interface between perception (MLLMs) and reasoning (LLMs). This allows the reasoning LLMs to be upgraded independently, boosting performance without the necessity to retrain the MLLMs or alignment. A detailed empirical analysis is provided in Sec. 4.3.
Figure 3: Comparison of the strategies for visual perception $O_{p}$.
Strategies for Visual Perception $O_{p}$.
We explore strategies to generate precise perceptual outputs for reasoning:
  • none: An empty set of perceptual outputs $O_{p}$, which serves as the control reference group.
  • cap (Lu et al., 2025): A holistic image caption $o_{\text{cap}}=\operatorname{MLLM}(I,P_{\text{cap}})$ with the template $P_{\text{cap}}$ in Fig. 13.
  • qcap (Gou et al., 2024): A query-relevant caption $o_{\text{qcap}}=\operatorname{MLLM}(I,P_{\text{qcap}}(q))$ with $P_{\text{qcap}}$ in Fig. 14.
  • sol: A tentative solution $o_{\text{sol}}=\operatorname{MLLM}(I,P_{\text{sol}}(q))$ with the template $P_{\text{sol}}$ in Fig. 15, which is “tentative” as it acts as contexts for LLMs instead of final answers.
To evaluate how different compositions of the perceptual output set $O_{p}$ affect performance, we perform an experiment on seven multi-modal reasoning benchmarks (details in Sec. 4.1 for details). In particular, we set $O_{p}$ to: (i) none; (ii) $\left\{o_{\text{cap}}\right\}$; (iii) $\left\{o_{\text{qcap}}\right\}$; (iv) $\left\{o_{\text{sol}}\right\}$; (v) $\left\{o_{\text{cap}},o_{\text{sol}}\right\}$; and (vi) $O_{p}=\left\{o_{\text{qcap}},o_{\text{sol}}\right\}$. The reasoning prompt $P_{r}$ (shown in Fig. 12) is designed to provide a unified structure for all the above cases. We use Qwen2.5-VL-7B for perception and adopt Qwen3-8B and R1-Distilled-7B (refered to as R1-7B) for reasoning. Figure 3 shows the average accuracies obtained. As can be seen,
  • Holistic captions outperform their query-relevant counterparts. This can be attributed to the MLLM’s extensive training on holistic image captioning tasks (Bai et al., 2025), whereas query-relevant captioning remains less optimized. However, with proper optimization (Sec. 3.2), query-relevant captioning can offer an advantage by extracting contextually relevant visual details.
  • Combining tentative responses and holistic captions achieves best results, delivering significant improvement (+7% w/ Qwen3-8B) over the original MLLM. This success is due to the complementary roles played by the caption and tentative response in reasoning. The caption provides the LLM with essential contexts for problem-solving, while the tentative response serves as a reference.
While Figure 3 shows that cap+sol performs best, we will adopt qcap+sol as the default in the sequel. The reason is empirical. Our findings in Section 4.2 reveal that qcap+sol outperforms cap+sol once VPO (to be introduced in Section 3.2) is applied. This indicates that the query-relevant approach, while less optimized initially, possesses greater potential.
3.2 Visual Perception Optimization (VPO)
Although the combination of caption and tentative solution (i.e., both cap+sol and qcap+sol) demonstrates superior results in Figure 3, they are not optimized for the correctness of the final reasoning outcome. In other words, the MLLM generates its perception outputs without any feedback on whether these outputs actually guide the reasoning LLM to the correct answer. To address this limitation, we introduce Visual Perception Optimization (VPO). As illustrated in Figure 4, VPO establishes a reinforcement learning feedback loop that fine-tunes the MLLM for better captioning, explicitly rewarding it based on the correctness of the final answer produced by the reasoning LLM.
Objective Design.